OpenStudy (anonymous):

Need Help BEST ANSWER! ! ! If the diameter of a circle is 25 inches, how long is the radius?

OpenStudy (anonymous):

r= c/2 25/2 = 12.5

OpenStudy (anonymous):

nvmmm!!!! i thought it wqas circumfrence~ sorry!

OpenStudy (anonymous):

Thank you so much your awesome and became your fan, do you know how to do radius

OpenStudy (anonymous):

I meant I'm try to find the radius lol any help will be great

OpenStudy (anonymous):

well radius is half of the circumfrence so just divide by 2 =]

OpenStudy (anonymous):

Ok thank you so I'm doing 25/5=5

OpenStudy (anonymous):

maybe u should ask somebody else, im a little bit rusty in this subject. idw you to get it wrong!!

OpenStudy (anonymous):

ok I think you right though

OpenStudy (anonymous):

Thank You !!!